ELECTRONIC SPECTROSCOPY OF THE A^{1}A′′-X^{1}A′ SYSTEM OF CHBr AND CDBr

\underline{M. DESELNICU}, C. TAO, C. MUKARAKATE AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53201-1881.

We report a comprehensive new study of the electronic spectroscopy of the A^{1}A′′-X^{1}A′ system of CHBr and CDBr in the region 450-750 nm. For each isotopomer, more than 30 cold bands in the progressions 2_{0}^{n}, 2_{0}^{n}3_{0}^{m} (m=1-3), 1_{0}^{1}2_{0}^{n},and 1_{0}^{1}2_{0}^{n}3_{0}^{m} (m=1-2) were observed, in addition to a number of hot bands. The spectra were measured under jet-cooled conditions using a pulsed discharge source, and rotationally analyzed to yield precise values for the band origins and rotational constants. The derived A^{1}A′′ state vibrational intervals are combined in Dixon plots with the results of Sears, Muckerman, and co-workers^{a} to derive barriers to linearity for the 2^{n}, 2^{n}3^{1}, and 2^{n}3^{2} progressions. The A^{1}A′′ C-H (C-D) stretching frequency was determined for the first time, and is in excellent agreement with the predictions of ab initio electronic structure theory.^{b} The observed ν_{3} dependence of the CH^{79}Br-CH^{81}Br isotope splitting in the A^{1}A′′ state is also in excellent agreement with theory.

aH.-G. Yu et al., J. Chem. Phys. 115, 5433 (2001).

bH.-G. Yu et al., J. Chem. Phys. 116, 1435 (2002).